A rectangular compost container is to be formed in a corner of a fenced yard, with 8 ft of chicken wire completing the other two sides of the rectangle. If the chicken wire is 3 ft high, what dimensions of the base will maximize the volume of the container?
Answer by oberobic(2304)

By theorem we know that to maximize the area of a quadrilateral we define a square.
Given 8 ft of fencing is to enclose 2 sides, then each side will be 4 ft.
The volume will be l*w*h = 4 * 4 * 3 = 48 cu ft.
You can experiment with other dimensions to ensure this is the maximum.